OpenStudy (anonymous):

Solve the equation shown for y. Your answer should be in y = mx + b form. 4x − 2y = 10

OpenStudy (across):

\[ 4x-2y=10\\ 2y=10-4x\\ y=-2x+5 \]EASY?